OpenStudy (anonymous):

Fan and medal! A rectangular sheet of paper is folded diagonally from K to N. Find the length of this diagonal.

OpenStudy (anonymous):

A.51 cm B.47 cm C.53 cm D.55 cm @kidrah69 @iGreen @Data_LG2

OpenStudy (igreen):

Use the pythagorean theorem again.

OpenStudy (igreen):

28^2 + 45^2 = c^2

OpenStudy (anonymous):

784+90=874

OpenStudy (igreen):

45^2 isn't 90

OpenStudy (anonymous):

sorry 2025+784=2809

OpenStudy (anonymous):

so 53

OpenStudy (igreen):

Yep, you got it.